\begin{thebibliography}{9}

\bibitem{app} P. Appell, {\it Sur les fonctions hyperg\'om\'etriques de plusieurs variables} In M\'emoir. Sci. Math. Paris: Gauthier-Villars, 1925.

\bibitem{appell} P. Appell and Kamp\'e de Feriet J, {\it Fonctions Hyperg\'eometriques et Hypersph\'eriques}, Polynome d'Hermite (Paris: Gautheir-Villars), 1926.

\bibitem{bailey} W. N. Bailey, {\it Appell's Hypergeometric Functions of Two Variables.}, Ch. 9 in Generalised Hypergeometric Series. Cambridge, England: Cambridge University Press, pp. 73-83 and 99-101, 1935.

\bibitem{beckert} U. Becker, N. Gr\"un and W. Scheid, {\it Evaluation of Appell functions in relativistic Coulomb-Dirac form factors}, J. Phys. A: Math. Gen. {\bf 20} (1987) 5447-5458.

\bibitem{Exton} H. Exton, {\it Multiple Hypergeometric Functions and Applications}, Halsted Press (Ellis Horwood Limited, Chichester), John Wiley and Sons, New York, 1976.

\bibitem{erdely1} A. Erd\'elyi, {\it Hypergeometric functions of two variables}, Acta Mathematica, 83 (1964) 131-164.

\bibitem{erdely2} A. Erd\'elyi, {\it Higher Transcendental Functions} Vol. 1, New York: McGraw-Hill, 1953.

\bibitem{taraso1} J. Fleischer, F. Jegerlehner and O. V. Tarasov, {\it A new hypergeometric next term representation of one-loop scalar integrals in $d$ dimensions}, Nuclear Physics B {\bf 672} (2003) 303-328.  

\bibitem{galue1} L. Galu\'e, {\it Una generalizaci\'on de la integral de Hubbell},
Rev. T\'ec. Ing. Univ. Zulia 14 (2), 153–160 (1991).

\bibitem{galue} L. Galu\'e, {\it Unification of generalized radiation field integrals involving confluent hypergeometric function}, Rad. Phys. Chem. 66 (2003) 269-273.

\bibitem{leubner} L. Galu\'e, S. L. Kalla and C. Leubner, {\it Algorithms for the approximation of the Hubbell rectangular source integral and its generalizations}, Radiat. Phys. Chem. 43 (1994) 497-502.

\bibitem{gabali} M. EL-Gabali and S. Kalla, {\it Some generalised radiation field integrals,} Computers Math. Applic 32 (1996) 121-128.

\bibitem{gabutti} B. Gabutti, S. L. Kalla and J. H. Hubbell, {\it  Some expansions
related to the Hubbell rectangular source integral.} J.
Comput. Appl. Math. 37, 273–285 (1991).

\bibitem{galue2} L. Galu\'e and V. Kiryakova, {\it Further results on a family of
generalized integrals,} Radiat. Phys. Chem. 43 (6), 573–579 (1994).

\bibitem{guseinev} I. I. Guseinov and B. N. Mamedov, {\it Calculation of the generalized Hubbell rectangular source integrals using binomial coefficients}, Appl. Math. Comput. 161 (2005) 285-292.

\bibitem{nasser2} Richard L. Hall, Nasser Saad, and Attila B von Keviczky, {\it Closed-form sums for some perturbation series involving associated Laguerre polynomials}, J. Phys. A: Math. Gen. 34, 11287-11300 (2001).

\bibitem{hubbell} J. H. Hubbell, R. L. Bach and J. C. Lamkin, {\it Radiation field from a rectangular source}, J. Res. Nat. Bur. Stand 64C (1960) 121-138. 

\bibitem{slkalla} S. Kalla, {\it The Hubbell rectangular source integral and
its generalizations}, Radiat. Phys. Chem. 41 (1993) 775–781.

\bibitem{KallaSL} S. L. Kalla, B. Al-Saqabi and S. Conde, {\it Some results related
to radiation field problems}, Hadronic J. 10 (1987) 221–230.

\bibitem{kallak} S. L. Kalla and H. G. Khajah, {\it Tau method approximation of some integrals related to radiation field problems}, Computers Math. Applic 33 (1997) 21-27. 

\bibitem{karule} S. L. Kalla, L. Galu\'e and V. Kiryakova {\it Some expansion related to a family of generalized radiation integrals}, Math. Balkanic  5 (1991) 190-202.

\bibitem{skalla} S. L. Kalla, C. Leubner, and B. Al-Saqabi, {\it Some results for the radiation flux in a rectangular straight duct}, Nucl. Sci. J. 28 (1991) 403-410.

\bibitem{Karule} E. Karule, {\it Integrals of two confluent hypergeometric functions},  J. Phys. A: Math. Gen. {\bf 23} (1990), 1969-1971.

\bibitem{landau} L. D. Landau and E. M. Lifshitz, {\it Quantum Mechanics}: {\it Non-Relativistic Theory}, Pergamon Press, London, 1981.

\bibitem{lee} J. L. Madajczyk and M. Trippenbach,  {\it Singular part of the hydrogen dipole matrix element}, {\it J. Phys. A}: {\it Math. Gen.} {\bf 22} (1989), 2369-2374.

\bibitem{opps} Sheldon B. Opps, Nasser Saad and H. M. Srivastava, {\it Some reduction and transformation formulas for the Appell hypergeometric function $F_2$\/}, J. Math. Anal. Appl. 302 (2005) 180-195.

\bibitem{Prudnikov} A.P. Prudnikov, Yu. A. Brychkov and O.I. Marichev, {\it Integrals and Series}, Vol. 3: {\it More Special Functions} (Translated from the 1983 Russian Edition by G.G. Gould), Gordon and Breach Science Publishers, New York, 1990

\bibitem{salman} M. Salman, {\it Generalized elliptic-type integrals and their representations}, Applied Mathematics and Computation {\bf 181} (2006) 1249–1256. 


\bibitem{nasser} Nasser Saad and Richard L. Hall, {\it Integrals containing confluent hypergeometric functions with applications to perturbed singular potentials}, {\it J. Phys. A}: {\it Math. Gen.} {\bf 36} (2003), 7771-7788.

\bibitem{nasser1} Nasser Saad and Richard L. Hall, {\it Closed-form sums for some perturbation series involving hypergeometric functions}, J. Phys. A.: Gen. Math. 35, 4105-4123 (2002)


\bibitem{slater} L. J. Slater, {\it Generalized Hypergeometric Functions}, Cambridge University Press, Cambridge, 1966.


\bibitem{sri} H. M. Srivastava and P. W. Karlsson, {\it Multiple Gaussian Hypergeometric Series}, Halsted Press (Ellis Horwood Limited, Chichester), John Wiley and Sons, New York, 1985.


\bibitem{tarasov} V. F. Tarasov,  W. Gordon's integral (1929) and its representations by means of Appell's functions $F_2$, $F_1$, and $F_3$, {\it J. Math. Phys.}, {\bf 44} (2003) 1449-1452.

\bibitem{trasov} V. F. Tarasov, {\it The generalizations of Slater's and Marvin's integrals and their representations
by means of Appell's functions $F_2(x,y)$}. Mod. Phys. Lett. B 8 (1994) 1417–1426.

\bibitem{trasov2} V. F. Tarasov, {\it Hypergeometric representation of the two-loop
equal mass sunrise diagram}, arXiv:hep-ph/0603227v1 (2006).


\end{thebibliography}

\documentclass[amsmath,amssymb]{revtex4}
\begin{document}
\title{Tables of the Appell Hypergeometric Functions $F_2$}

\author{Jonathan Murley}%
\author{Nasser Saad}
\email{Jmurley@stu1.cs.upei.ca}
\email{nsaad@upei.ca}
\affiliation{Department of Mathematics and Statistics, University of Prince Edward Island 
Charlottetown, Prince Edward Island C1A 4P3, Canada}%


\def\dbox#1{\hbox{\vrule  
                        \vbox{\hrule \vskip #1
                             \hbox{\hskip #1
                                 \vbox{\hsize=#1}%
                              \hskip #1}%
                         \vskip #1 \hrule}%
                      \vrule}}
\def\qed{\hfill \dbox{0.05true in}}  
\def\square{\dbox{0.02true in}} 
\begin{abstract}
\noindent{\bf Abstract:} The generalized hypergeometric function $_qF_p$ is a power series in which the ratio of successive terms is a rational function of the summation index.  The Gaussian hypergeometric functions $_2F_1$ and $_3F_2$ are most common special cases of the generalized hypergeometric function $_qF_p$. The Appell hypergeometric functions $F_q$, $q=1,2,3,4$ are  product of two hypergeometric functions $_2F_1$  that appear in many areas of mathematical physics. Here, we are interested in the Appell hypergeometric function $F_2$ which is known to have a double integral representation. As demonstrated by Opps, Saad, and Srivastava (J. Math. Anal. Appl. 302 (2005) 180-195), the double integral representation of $F_2$ can be reduced to a single integral that can be easily evaluated for certain values of the parameters in terms of $_2F_1$ and $_3F_2$. Using many of the reduction formulas of $_2F_1$ and $_3F_2$ and the representation of $F_2$ in terms of a single integral, we have begun to tabulate new reduction formulas for $F_2$.
\end{abstract}
\vskip0.1 true in
\maketitle
\noindent{\bf PACS:} {Primary 33C65, 33C05, 33D15, 33D60; Secondary 33B15, 33C20, 33D90.}
   \vskip0.1true in                        
\noindent{\bf keywords:} {~Multiple hypergeometric functions, Appell series, Generalized hypergeometric function, Gauss hypergeometric function,  Clausen hypergeometric function,  Reduction and transformation formulas, Radiation field, Hubbell integral.
}


\section{Introduction}	
\noindent  Appell hypergeometric functions $F_D,D=1,2,3,4$  play an important role in mathematical physics (\cite{app}-\cite{bailey}, \cite{Exton}, \cite{erdely1}-\cite{erdely2}, \cite{slater}, \cite{sri}). In particular, the Appell hypergeometric series $F_2$ arises frequently in various physical and chemical applications (\cite{beckert}, \cite{taraso1} to \cite{opps}, \cite{salman} to \cite{trasov2} ). The exact solution of number of problems in quantum mechanics has been given \cite{landau} in terms of Appell's function $F_2$.
It is defined by [\cite{slater}, p. 211, Equation (8.1.4)]: 
\begin{equation}\label{eq1}
F_2(\sigma,\alpha_1,\alpha_2;\beta_1,\beta_2;x,y)=\sum\limits_{m=0}^\infty\sum\limits_{n=0}^\infty {(\sigma)_{m+n}(\alpha_1)_m(\alpha_2)_n\over (\beta_1)_m(\beta_2)_n} \, {x^m \over m!} \, {y^n \over n!},
\end{equation}
for $|x| + |y| <1; \, \beta_j \in {\Bbb C} \backslash {\Bbb Z}_0^-; \, {\Bbb Z}_0^-:=\lbrace 0,-1,-2, \ldots \rbrace,$ and
$(\lambda)_k$ denotes  the Pochhammer symbol defined, in terms of Gamma functions, by
$$(\lambda)_k:= {\Gamma(\lambda+k)\over\Gamma(\lambda)}=\left\{ \begin{array}{ll}
 1 &\mbox{ if\quad $\left( k=0; \, \lambda \in {\Bbb C} \backslash \lbrace 0 \rbrace \right) $} \\
  \lambda(\lambda+1)(\lambda+2)\dots(\lambda+k-1) & \mbox{ if\quad $\left( k \in {\Bbb N}; \, \lambda \in {\Bbb C} \right)$}
       \end{array} \right.
$$
where $\Bbb N$ being the set of {\it positive} integers. Further, it also has the following double integral representation [\cite{slater}, p. 214, Equation (8.2.3)]:
\begin{eqnarray}\label{eq2}
F_2(\sigma,\alpha_1,\alpha_2;\beta_1,\beta_2;x,y)&=&{\Gamma(\beta_1)\Gamma(\beta_2)\over \Gamma(\alpha_1)\Gamma(\alpha_2)\Gamma(\beta_1-\alpha_1)\Gamma(\beta_2-\alpha_2)}\int_0^1
\int_0^1u^{\alpha_1-1}\tau^{\alpha_2-1}\nonumber\\
&~&\times(1-u)^{\beta_1-\alpha_1-1}(1-\tau)^{\beta_2-\alpha_2-1}(1-xu-y\tau)^{-\sigma}du \, d\tau
\end{eqnarray}
where ${\frak R}\left( \beta_j \right) > {\frak R} \left( \alpha_j \right) >0, \, j=1,2$ and $|x| +|y| <1$. Recently, Opps et al.  \cite{opps} used the Euler's integral representation of the Gauss hypergeometric function [\cite{slater}, p. 20, Equation (1.6.6)] 
\begin{equation}\label{eq3}
{}_2F_1\bigg({a,b \atop c} \bigg|z \bigg)={\Gamma(c)\over \Gamma(b)\Gamma(c-b)}
\int_0^1\tau^{b-1}(1-\tau)^{c-b-1}(1-z\tau)^{-a}d\tau 
\end{equation}
to reduce the double integral representation (\ref{eq2}) into a single integral in term of $_2F_1$, 
\begin{equation}\label{eq4}
F_2(\sigma;\alpha_1,\alpha_2;\beta_1,\beta_2;x,y)
={\Gamma(\beta_1)\over \Gamma(\alpha_1)\Gamma(\beta_1-\alpha_1)}
\int_0^1 {u^{\alpha_1-1}(1-u)^{\beta_1-\alpha_1-1}\over (1-xu)^{\sigma}}{}_2F_1\bigg({\sigma,\alpha_2 \atop \beta_2}\bigg|{y \over 1-xu}\bigg)~du
\end{equation}
for ${\frak R}\left( \beta_1 \right) > {\frak R}\left( \alpha_1 \right) >0$, and $|x| + |y| <1.$ Using some properties of $_2F_1$, they prove the following theorem [\cite{opps}, Theorem 1]:
\vskip0.1true in
\noindent {\bf Theorem 1.} {\it For $|x|+|y|<1$, the Appell hypergeometric function $F_2$ is given by
\begin{eqnarray}\label{eq5}
F_2(a+1;\alpha_1,1;\beta_1,2;x,y)&= & -{1\over ay}\ 
{}_2F_1\bigg({a,\alpha_1 \atop \beta_1} \bigg|x\bigg) +{1\over ay(1-y)^{a}}\ {}_2F_1\bigg({a,\alpha_1 \atop \beta_1} \bigg|{x\over 1-y}\bigg)
\end{eqnarray}
where $a \neq 0; \, \alpha_1 \in {\Bbb C}; \, \beta_1 \in {\Bbb C} \backslash {\Bbb Z}_0^-$ and
\begin{eqnarray}\label{eq6}
F_2(1;\alpha_1,1;\beta_1,2;x,y)&= & \,
{\alpha_1 \over \beta_1y}\bigg[{x\over 1-y}\ 
{}_3F_2\bigg( {\alpha_1+1,1,1 \atop \beta_1+1,2}\bigg|{x\over 1-y}\bigg)- x\ {}_3F_2\bigg({\alpha_1+1,1,1 \atop \beta_1+1,2}\bigg|x\bigg)\bigg]-{\ln(1-y)\over y}\nonumber\\
\end{eqnarray}
where $\alpha_1 \in {\Bbb C}; \, \beta_1 \in {\Bbb C} \backslash {\Bbb Z}_0^-.$}
\noindent  The present work is devoted to compute $F_2(\sigma;\alpha_1,1;\beta_1,2;x,y)$ explicitly for different values $\sigma,\alpha_1,\beta_1$ of the function parameters using (\ref{eq5}) and (\ref{eq6}). This is mostly done using many of the reduction formulas of $_2F_1$ and $_3F_2$ listed in \cite{Prudnikov} and other sources of special functions (\cite{Exton}, \cite{erdely1}-\cite{erdely2}, \cite{slater}, \cite{sri}), we begun here to tabulate reduction and transformation formulas for $F_2$. First, in Tables I and II, we give the corrections to some formulas misprinted in the classical monograph by Prudnikov et al. \cite{Prudnikov}.
\begin{table}[h1]
\begin{center}
\caption{Correction to some formulas for ${}_2F_1$ reported in the classical work of Prudnikov et al \cite{Prudnikov}. }
\vspace{0.1in}
\begin{tabular*}{1.0\textwidth}{@{\extracolsep{\fill}} | l || l || l || l | }
  \hline
\textbf{$\sigma$} &  $\alpha$    & $\beta$  & $_2F_1(\sigma,\alpha;\beta;z)$ \\
  \hline\hline
$\frac{5}{2}$ & $4$ & $1$ & $\frac{1}{16}(16+72z+18z^2-z^3)(1-z)^{-\frac{11}{2}}$ \\
\hline
 \hline
$\frac{4}{5}$ & $1$ & $\frac{14}{5}$ & $\frac{9}{5x^5}-\frac{9}{25x^9}(1-x^5)\big[\ln(1-x^5)-5\ln(1-x)-\sqrt{5}\ln\frac{1-2^{-1}(\sqrt{5}-1)x+x^2}{1+2^{-1}(\sqrt{5}+1)x+x^2}-$\\
~&~&~&$\quad 2(10+2\sqrt{5})^\frac{1}{2}\arctan\frac{(10+2\sqrt{5})^\frac{1}{2}x}{4-(\sqrt{5}-1)x}-2(10-2\sqrt{5})^\frac{1}{2}\arctan\frac{(10-2\sqrt{5})^\frac{1}{2}x}{4+(\sqrt{5}+1)x}\big]\qquad\qquad\qquad [x=z^\frac{1}{5}]$ \\
\hline
 \hline
$\frac{5}{6}$ & $1$ & $\frac{17}{5}$ & $\frac{11}{6x^6}+\frac{55}{36x^{11}}(1-x^6)(\ln\frac{1-x}{1+x}+\frac{1}{2}\ln\frac{1-x+x^2}{1+x+x^2}+\sqrt{3}\arctan\frac{3^\frac{1}{2}x}{1-x^2}\big]$ \\
\hline
 \hline
$1$ & $\frac{7}{2}$ & $\frac{9}{2}$ & -$\frac{7}{15z^3}\big[15+15z+3z^2-\frac{15\tanh^{-1}\sqrt{z}}{\sqrt{z}}\big]$ \\
\hline
 \hline
$1$ & $b$ & $b-m$ & $\frac{b-m-1}{b-1}\sum_{k=0}^{m-1}\frac{(-m)_k}{(2-b)_k}(1-z)^{-k-1}-\frac{m!}{(1-b)_m}(z-1)^{-m-1}$ \\
\hline
 \hline
$-\frac{n}{2}$ & $\frac{1-n}{2}$ & $1-n$ & $2^{-n}(1+\sqrt{1-z})^n,\qquad\qquad(n\neq{1, 2})$ \\
\hline
 \hline
\end{tabular*}
\end{center}
\end{table}
\begin{table}[h2]
\begin{center}
\caption{Correction to some formulas for $_3F_2$ reported in the classical work of Prudnikov et al \cite{Prudnikov}. }
\vspace{0.1in}
\begin{tabular*}{1.0\textwidth}{@{\extracolsep{\fill}} | l | l | l || l| l|| l | }
  \hline
\textbf{$a_1$} &  $a_2$    & $a_3$  & $b_1$& $b_2$  & $_3F_2(a_1,a_2,a_3;b_1,b_2;z)$ \\
  \hline\hline
$\frac{1}{4}$ &  $1$    & $1$  & $\frac{5}{4}$& $2$  & $\frac{1}{3z}\big[\ln(1-z)+z^\frac{3}{4}\big(\ln\frac{1+z^\frac{1}{4}}{1-z^\frac{1}{4}}+2\arctan{z^\frac{1}{4}}\big)\big]$\\
\hline
  \hline
\end{tabular*}
\end{center}
\end{table}
\section{Special values of the Appell Hypergeometric Functions $F_2$}

\noindent In the next, we tabulate the explicit computations of  $F_2(\sigma;\alpha_1,1;\beta_1,2;x,y)$ for different values of the function parameters $\sigma,\alpha_1$, and $\beta_1$. Since the role of $\alpha_1, \beta_1,x$ and $\alpha_2, \beta_2,y$ in $F_2(\sigma;\alpha_1,\alpha_2;\beta_1,\beta_2;x,y)$ are interchanged, similar tables can be obtain for  $F_2(\sigma;1,\alpha_2;2,\beta_2;x,y)$. Even-though the table is given only for $\alpha_2=1$ and $\beta_2=2$, the table can be used for a wider variety of cases. This can be notice from the following 
properties.
\begin{enumerate}
  \item $F_2(\sigma,\alpha_1,\alpha_2;\beta_1,\beta_2;x,y)=F_2(\sigma,\alpha_2,\alpha_1;\beta_2,\beta_1;y,x)$.
  \item $F_2(\sigma,\alpha_1,\alpha_2;\beta_1,\beta_2;x,y)=(1-x)^{-\sigma}F_2(\sigma,\beta_1-\alpha_1,\alpha_2;\beta_1,\beta_2;{x\over x-1},{y\over 1-x})=(1-x-y)^{-\sigma}F_2(\sigma,\beta_1-\alpha_1,\beta_2-\alpha_2;\beta_1,\beta_2;{x\over x+y-1},{y\over x+y-1})$.
  \item $F_1(\alpha;\beta,\beta',\gamma;x,y)=\left({x\over y}\right)^{\beta'}F_2(\beta+\beta';\alpha,\beta';\gamma,\beta+\beta';x,1-{x\over y})=\left({y\over x}\right)^{\beta}F_2(\beta+\beta';\alpha,\beta;\gamma,\beta+\beta';y,1-{y\over x}).$
 \end{enumerate}  
\noindent Further, the analytic expressions presented in the following tables can be useful in many applications of mathematical physics including the computation of the generalized Hubbell rectangular source integrals, elliptic integrals, and the radiation fields (\cite{beckert}, \cite{taraso1} to \cite{opps}, \cite{salman} to \cite{trasov2}).  

\begin{center}

\end{center}

\section{Representations of Appell hypergeometric function $F_2$}

\begin{center}
%
\end{center}

\bigskip
\section*{Acknowledgments}
\medskip
\noindent The present study was supported, in part, by the {\it Natural Sciences and Engineering Research Council of Canada} under Grant GP249507 (NS).  
\medskip



\end{document}